The box-and-whisker plots show the distribution of the masses (in grams) of eggs sold in Shop A and Shop B.

  1. (i)

    There are 25 eggs with masses of more than 59 g sold in Shop A.

    Find the total number of eggs sold in Shop A.

    Answer ... eggs

  2. (ii)

    Make a comment comparing the averages and a comment comparing the distribution of the masses of the eggs sold in Shop A and Shop B.

    Use figures to support your answers.

    1. ...

    ...

    1. ...

    ...

  3. (8)
    1. (b)

      In a sample of 80 eggs, 4 are cracked.

      1. (i)

        One egg is selected from the sample at random.

        Find the probability that the egg is cracked.

        Answer ...

      2. (ii)

        Two eggs are selected from the sample at random.

        Find the probability that both eggs are cracked.

        Answer ...

      3. (iii)

        Three eggs are selected from the sample at random.

        Find the probability that at least one egg is cracked.

        Answer ...

Work through it with AI →

It won't just hand you the answer — it asks where you're stuck.How the checker works

Worked solution

8 (a)(i)

1. Relate masses above the upper quartile to one quarter of the total number of eggs.

For Shop A, Q3=59Q_3 = 59 g.

So the eggs with masses more than 5959 g make up 25%25\% of the total.

25% of total=2525\% \text{ of total} = 25

Total=25÷0.25=100\text{Total} = 25 \div 0.25 = 100

Answer

100100 eggs

8 (a)(ii)

1. Compare the averages using the medians from the box-and-whisker plots.

For a box-and-whisker plot, the average is usually compared using the median.

Shop A median =54=54 g.

Shop B median =50=50 g.

Since 54>5054>50, Shop A has the higher average mass.

2. Compare the distributions using a measure of spread from the box-and-whisker plots.

Use the interquartile range to compare spread.

Shop A:
IQR=Q3−Q1=59−48=11\text{IQR}=Q_3-Q_1=59-48=11 g.

Shop B:
IQR=Q3−Q1=58−46=12\text{IQR}=Q_3-Q_1=58-46=12 g.

Since 12>1112>11, Shop B has a slightly greater spread in the middle 50%50\% of the masses.

Answer
  1. The average mass of eggs in Shop A is higher than in Shop B, since the median for Shop A is 5454 g while the median for Shop B is 5050 g.

  2. The masses of eggs in Shop B are more spread out in the middle 50%50\% of the data, since its interquartile range is 58−46=1258-46=12 g, while Shop A's interquartile range is 59−48=1159-48=11 g.

8 (a)(8)(b)(i)

1. Find the probability that the selected egg is cracked.

There are 44 cracked eggs out of 8080 eggs.

P(cracked)=480=120P(\text{cracked})=\frac{4}{80}=\frac{1}{20}

Answer

120\dfrac{1}{20}

8 (a)(8)(b)(ii)

1. Find the probability that the first egg selected is cracked.

There are 44 cracked eggs out of 8080 eggs.

P(first cracked)=480P(\text{first cracked})=\frac{4}{80}

2. Account for no replacement and multiply the probabilities.

After selecting one cracked egg, there are 33 cracked eggs left out of 7979 eggs.

P(both cracked)=480×379=126320=31580P(\text{both cracked})=\frac{4}{80}\times\frac{3}{79}=\frac{12}{6320}=\frac{3}{1580}

Answer

1158\dfrac{1}{158}

8 (a)(8)(b)(iii)

1. Find the probability that none of the three selected eggs is cracked.

There are 80−4=7680-4=76 uncracked eggs.

P(none cracked)=7680×7579×7478P(\text{none cracked})=\frac{76}{80}\times\frac{75}{79}\times\frac{74}{78}

2. Use the complement to find the probability that at least one egg is cracked.

P(at least one cracked)=1−7680×7579×7478P(\text{at least one cracked})=1-\frac{76}{80}\times\frac{75}{79}\times\frac{74}{78}

=1−56157900=1-\frac{5615}{7900}

=22857900=229790=\frac{2285}{7900}=\frac{229}{790}

Answer

229790\dfrac{229}{790}

AI analysis — may contain mistakes.

Worked this one out on paper? Snap your steps and Gigle's AI math checker will circle the step where it went wrong.

More questions on this topic

  1. 1The masses of two batches of 400 steel bars produced by factory A and B were measured and recorded. The cumulative frequ…
  2. 2The stem-and-leaf diagram shows the test scores of the boys and girls from a particular class. [Stem-and-leaf diagram pr…
  3. 3The cumulative frequency graph shows distribution of the waiting times before being served, in minutes, of 60 customers…
  4. 4The waiting time, t in minutes, for 40 customers at two phone service shops, Shop A and Shop B, are as follows. Calculat…
  5. 5The temperatures of eighteen girls in a class on a particular day are shown in the stem-and-leaf diagram. (Stem-and-leaf…